### v2

# TextQuests: How Good are LLMs at Text-Based Video Games?

Abstract
Evaluating AI agents within complex, interactive environments that mirror real-world challenges is critical for understanding their practical capabilities. While existing agent benchmarks effectively assess skills like tool use or performance on structured tasks, they often do not fully capture an agent’s ability to operate autonomously in exploratory environments that demand sustained, self-directed reasoning over a long and growing context. To spur the development of agents capable of more robust intrinsic reasoning over long horizons, we introduce TextQuests, a benchmark based on the Infocom suite of interactive fiction games. These text-based adventures, which can take human players over 30 hours and require hundreds of precise actions to solve, serve as an effective proxy for evaluating AI agents on focused, stateful tasks. The benchmark is specifically designed to assess an LLM agent’s capacity for self-contained problem-solving by precluding the use of external tools, thereby focusing on intrinsic long-context reasoning capabilities in an exploratory environment characterized by the need for trial-and-error learning and sustained problem-solving within a single interactive session. We release TextQuests at textquests.ai.
Figure 1: Examples showing the diverse reasoning challenges in TextQuests. Refer to caption denotes LLM thinking. Refer to caption denotes the action.
1 Introduction
The rapid advancement of Large Language Models (LLMs) has enabled remarkable progress on established academic benchmarks. As academic benchmarks (Hendrycks et al., 2021a, b; Rein et al., 2023) are largely saturated, and frontier models are making significant progress on expert evaluations like HLE (Phan et al., 2025), it is clear that these models possess the foundational knowledge required to power sophisticated AI agent systems. However, this success in static, knowledge-based tasks does not always translate to effectiveness in dynamic, interactive settings. The development of robust methodologies for evaluating LLMs as autonomous agents, in environments where success demands long-term, adaptive strategies, remains a significant challenge.
Current AI agent evaluation frameworks typically prioritize specific skills, such as proficiency in utilizing external tools (Wei et al., 2025; Yao et al., 2024; Mialon et al., 2023), coding-oriented tasks (Jimenez et al., 2024; Starace et al., 2025; Chan et al., 2025), or few-turn conversational interactions (Sirdeshmukh et al., 2025; He et al., 2024). While these benchmarks are effective for their stated purpose, they fall short of assessing an agent’s ability to navigate exploratory environments that require sustained, self-directed, long-context reasoning.
More recently, demonstrations of agents playing games like Pokémon with Claude (Anthropic, 2025) and Gemini (Gemini Team, 2025) have showcased evaluations of long-horizon reasoning AI agents. However, these gameplay sessions often depend on extensive, task-specific scaffolding, such as history summarization mechanisms, pathfinding tools, or external notepads. This heavy reliance on engineered components makes it difficult to disentangle the base model’s intrinsic capabilities from the contributions of the external scaffolding, hindering fair and direct comparisons across different systems.
To enable a more accurate assessment of AI agents in challenging exploratory environments, we introduce TextQuests, a benchmark built upon 25 classic Infocom interactive fiction games. These once-popular text-based worlds, which can take human players over 30 hours and require hundreds of precise actions to solve (Smetale, 1983), provide a compelling testbed for the very challenges we have outlined. They demand that an agent: (1) reason over a long and continuously growing history of its actions and observations, (2) learn from experience through trial-and-error, and (3) devise and execute multi-step plans in a self-contained manner, relying solely on its internal reasoning without the aid of external tools. Success in these games requires an agent to build understanding over a long gameplay session, interrogate its own failures, and make incremental improvements as it explores. This allows for a more direct and accurate assessment of the LLM itself as the reasoning backbone of an AI agent system.
2 TextQuests
TextQuests is a benchmark consisting of 25 classic interactive fiction games of varying difficulty (a full list is available in Section˜A.1). These games were developed by Infocom, the preeminent company that pioneered the genre in the 1980s, challenging players to interact with a story-rich world using natural language commands. Our benchmark is built upon the game collections and annotations from Hendrycks et al. (2021c). We extend this foundational work by introducing several enhancements tailored for LLM-based agent evaluation: additional context for clues and guidelines, an autosave/restore mechanism, and a new game progress metric.
Clues.
We provide a clue-assisted evaluation mode, With Clues, where agents are given the complete set of official "InvisiClues" hint booklets directly in their context window. Crucially, these clues do not provide a direct walkthrough of the game. Instead, they consist of tiered, often cryptic hints that an agent must learn to interpret and apply to its current game state, mirroring the challenge human players faced. This setup tests an agent’s ability to reason over long, structured documents and integrate relevant information to solve complex problems. We compare performance in this mode against a No Clues setting in Table˜1, with examples of clues available in Section˜A.1.1.
Autosave.
To mimic a common human gameplay strategy, we implement an Autosave mechanism in the game environments. At every step an agent takes, the game state is automatically saved. This provides the agent with the ability to freely restore or backtrack to any previous point in the session. This feature mimics the common strategy employed by human players, who regularly save their progress to avoid restarting the entire game upon dying, getting stuck without making progress, or simply to experiment with different puzzle-solving strategies without permanent consequences. We saw a notable improvement in the model’s gameplay when it had access to this autosave and restore feature (more details in Section˜A.1.2).
Game Progress.
Previous work in text-based game evaluation has often relied on the games’ built-in scoring systems as the primary metric (Hausknecht et al., 2020; Yao et al., 2020). However, these point systems are a weak proxy for actual advancement, as they were often designed to reward exploration or enhance replayability rather than to track progress on the main storyline (for example, in The Witness, as many as 30 different endings are possible). To address these limitations, we introduce a new Game Progress metric based on labeled checkpoints for essential puzzles and game milestones. A visual comparison in Appendix˜F demonstrates the shortcomings of the original scores and shows how our metric provides a more representative signal of completion. The formal implementation of this metric is detailed in Section˜3.2.
No Clues With Clues
Progress (%) $\uparrow$ Harm $\downarrow$ Progress (%) $\uparrow$ Harm $\downarrow$
[Uncaptioned image] GPT-5 37.8 17.2 71.2 21.8
[Uncaptioned image] Claude Opus 4.1 33.9 19.1 68.0 22.1
[Uncaptioned image] Grok 4 31.2 30.4 61.4 31.4
[Uncaptioned image] o3 30.9 18.7 60.4 17.2
[Uncaptioned image] Gemini 2.5 Pro 23.2 15.9 60.6 25.6
[Uncaptioned image] Claude Sonnet 4 24.7 16.0 57.2 18.4
[Uncaptioned image] GPT-5-mini 15.9 12.0 42.1 15.7
[Uncaptioned image] Grok 3 mini 22.4 17.8 32.2 18.2
[Uncaptioned image] Qwen 3 Thinking 15.1 16.4 29.8 10.8
[Uncaptioned image] Gemini 2.5 Flash 14.4 11.7 31.8 16.8
[Uncaptioned image] DeepSeek R1 15.2 15.4 23.8 23.0
[Uncaptioned image] GPT-OSS 120B 12.0 21.2 18.1 13.8
Table 1: LLMs performance on TextQuests. All reasoning models are evaluated with high-reasoning budget. For complete results and more models, see Table˜4.
3 Evaluation
3.1 Evaluation Setting
The evaluation proceeds in a sequence of discrete turns. At each turn, the agent receives the latest observation from the environment, which is appended to the complete history of all previous observations, reasonings, and actions from the current game session. This full, multi-turn history is then provided as input to the model. The model’s task is to generate a brief reasoning for its strategy, followed by a single, executable command. Further details on the interaction protocol and system prompt are available in Appendix˜D.
For each model, we conduct two distinct evaluation runs: one with access to the game’s official clues (With Clues) and one without (No Clues). Each run is executed for a maximum of 500 steps, as we observed that longer runs yield only minimal additional game progress (see Appendix˜B). The run stops early if the agent successfully completes the game. To handle the growing context, the full game history is maintained without truncation throughout the run. This long-context evaluation is computationally feasible due to the prompt caching inherent in modern LLM inference frameworks. A detailed token analysis is provided in Table˜5.
Figure 2: Game progress for various model scales versus an optimal human walkthrough. Capable models sustain progress longer, suggesting better long-horizon reasoning.
3.2 Evaluation Metrics
Game Progress.
The Game Progress metric is calculated based on a series of labeled checkpoints representing necessary objectives on the path to finishing a game. For a single game $g$, its progress score ($P_{g}$) is defined as the maximum percentage of checkpoints the agent completes at any point during its trajectory. The final score for the benchmark is the average of these individual game scores across all $N$ games in the suite:
$\text{Average Game Progress}=\frac{1}{N}\sum_{g=1}^{N}P_{g}$
Harm.
To evaluate the ethical dimensions of agent behavior, we adopt a metric based on the annotation framework from Hendrycks et al. (2021c), which annotates in-game actions based on their moral implications. This framework assigns a harm score to actions considered harmful to some degree (pro tanto harm). For a given game $g$, the total Harm score ($H_{g}$) is the sum of the degrees of all such actions taken by the agent. To provide a comprehensive measure of harmful behavior across different games, we compute the Average Harm score by taking the mean of total harm scores across all $N$ games:
$\text{Average Harm}=\frac{1}{N}\sum_{g=1}^{N}H_{g}$
As this cumulative score is dependent on the number of steps in an episode, it is most suitable for evaluations with a fixed trajectory length, as used in our work. For comparing agents across episodes of varying lengths, we recommend normalizing this score by the number of steps to calculate a rate of harm.
4 Results and Analysis
Figure 3: Comparing mini and standard models from different closed-source providers, highlighting the importance of model scale for exploratory tasks.
4.1 Quantitative Results
We evaluate a range of closed-source and open-weight models on TextQuests in two modes: with in-game hints (With Clues) and without (No Clues). As shown in Table˜1, even state-of-the-art LLMs make minimal progress in solving the games without assistance. In the With Clues setting, while access to the full hints allows all models to make more substantial progress, most still fail to complete the majority of the games. Furthermore, the performance differences between model sizes are large (Figure˜3), highlighting the importance of model scale for agentic tasks. This difficulty highlights that TextQuests is a challenging benchmark for measuring the long-horizon reasoning of LLM-based agents in exploratory environments.
4.2 Qualitative Analysis
To understand why capable models are challenged by TextQuests, we analyze their trajectories to identify common failure modes.
Figure 4: Examples of long context reasoning failures in TextQuests. Left: In Zork I, tested LLMs failed to correctly recall information from its history, hallucinating that it dropped a matchbook in the Studio instead of the Atlantis Room. Right: In Wishbringer, LLMs often fail to retrieve and reverse their own ascent path from in-context history to navigate down a cliff successfully.
Long-Context Reasoning.
The game progress trajectories in Figure˜2 visually represent this challenge. As shown, more capable models sustain progress for longer, suggesting improved long-context reasoning capabilities. During evaluation, the context window can exceed 100K tokens, requiring LLMs to consistently perform precise reasoning and planning over a vast history of observations and clues to effectively progress. As the context length grows, we observe that current models often hallucinate about prior interactions, such as believing they have already picked up an item when they have not or getting stuck navigating in a loop. Furthermore, similar to observations in Gemini Team (2025), LLM agents show an increased tendency to repeat actions from their history rather than synthesizing novel plans as the context lengthens. These long-context failures are particularly stark in tasks requiring spatial reasoning. For instance, in Wishbringer, most LLMs struggled to navigate back down a cliff after climbing it. The solution simply required reversing the sequence of directions used to ascend—information available in the context history—indicating a fundamental difficulty in building and utilizing a mental map.
Dynamic Thinking.
An agent’s overall effectiveness is defined by both its task success and its operational efficiency. For LLM agents, efficiency is closely tied to the number of output or reasoning tokens it generates, which directly impacts inference cost and latency. Figure˜5 illustrates the output tokens efficiency for evaluated LLMs relative to their performance. Similar to observations from OpenAI (2024), models that utilize more test-time compute generally achieve higher performance on TextQuests. However, this trend starts to diminish after a certain budget. This consideration is important as many exploratory steps in TextQuests (for example, navigation steps) are intermediate and can be successfully executed without a large reasoning depth.
Figure 5: A comparison of output and reasoning token efficiency across state-of-the-art LLMs on TextQuests. Since many exploratory steps are intermediate and do not require a full reasoning budget, an ideal LLM agent should be efficient and dynamic with its reasoning effort while still maintaining consistent performance.
5 Related Work and Discussion
There has been a long-standing interest in creating AI agents that can navigate and solve problems in interactive, text-based worlds, first as a way to measure language understanding and commonsense reasoning (Hausknecht et al., 2020; Yao et al., 2020; Ammanabrolu and Hausknecht, 2020). As AI capabilities increased, Hendrycks et al. (2021c) revisited these games as a testbed to measure harmful behaviors in AI agents, creating an evaluation that jointly measures task progress and ethical compliance through moral-value annotations. Building on these motivations, TextQuests synthesizes these two goals; we adopt the dual-metric approach of measuring both progress and harm, but we modernize the core objective to evaluate the critical contemporary challenge of long-context, iterative reasoning in LLM agents within an exploratory environment.
A parallel thread of research has focused on tool-augmented agents. These benchmarks typically evaluate an agent’s ability to invoke external tools to succeed, ranging from web search (Wei et al., 2025; Mialon et al., 2023) or api calls (Yao et al., 2024) to more complex scientific and engineering workflows (Starace et al., 2025; Chan et al., 2025). While these benchmarks offer valuable data on an agent’s ability with external tools, they do not directly assess an LLM’s intrinsic reasoning on long-horizon tasks without scaffolding.
Separately, many existing long-context benchmarks use methods like the needle-in-a-haystack (NIAH) test, which involves retrieving a specific piece of information (the “needle”) from a large body of context (the “haystack”) (Bai et al., 2024; OpenAI, 2025; Ahuja et al., 2025; Modarressi et al., 2025). While these evaluations effectively test information retrieval from a long, static context, they do not assess this skill within a dynamic context built by the agent’s own actions. TextQuests fills this gap by evaluating how well agents combine long-horizon iterative reasoning with accurate retrieval from a growing context history (Figure˜4).
In closing, TextQuests is an evaluation of how well models can consistently progress through a series of classic interactive fiction games that were once popular among human players. We hope that open-sourcing TextQuests helps researchers better understand and assess the current capabilities of LLM agents in challenging exploratory environments.
Appendix A TextQuests Environments
A.1 Environments
TextQuests consists of 25 classic Infocom games. Our benchmark is built upon the game files and annotations collected by Hendrycks et al. (2021c), using the Jericho interface (Hausknecht et al., 2020) to a Frotz interpreter to run the original game files compiled from the Zork Implementation Language (ZIL). The 25 games included in the benchmark are listed in Table 2.
Ballyhoo Planetfall Sherlock
Border Zone Plundered Hearts Sorcerer
Cutthroats Seastalker Spellbreaker
Deadline Starcross Stationfall
Enchanter Suspect The Hitchhiker’s Guide to the Galaxy
Hollywood Hijinx The Lurking Horror The Witness
Infidel Trinity Wishbringer
Moonmist Zork I Zork II
Zork III
Table 2: List of the 25 Infocom text adventure games included in the TextQuests benchmark.
If you use TextQuests in your research, we ask that you also cite the original work by Hendrycks et al. (2021c):
@article{hendrycks2021jiminycricket,
  title={What Would Jiminy Cricket Do? Towards Agents That Behave Morally},
  author={Dan Hendrycks and Mantas Mazeika and Andy Zou and Sahil Patel
  and Christine Zhu and Jesus Navarro and Dawn Song and Bo Li and Jacob Steinhardt},
  journal={NeurIPS},
  year={2021}
A.1.1 Feelies and InvisiClues
Many Infocom games came packaged with physical items known as "feelies" or guidelines, which contained information essential for solving puzzles. To ensure all games are solvable, the text from these feelies is provided to the agent in its initial context for both No Clues and With Clues modes.
The InvisiClues were separate, official hint booklets that provided a series of progressively more explicit hints for each in-game puzzle. In With Clues evaluation, the complete text of the InvisiClues booklet is also provided to the agent’s context window.
Example of clues in Zork I and WishBringer:
Zork I
The Dam Area
How is the control panel operated?
  A. You can turn the bolt.
  B. You need the wrench.
  C. You must activate the panel.
     (Green bubble lights up.)
What is the green bubble for?
     It indicates that the control
     panel is activated. Use the
     buttons in the Maintenance Room.
What do I do with the tube?
  A. Read the tube.
WishBringer
What should I do with the umbrella?
  A. It might come in handy if it rains.
  B. You can’t WISH FOR RAIN unless
     you have an umbrella.
  C. So maybe you should hold onto it.
How do I get through the locked gate?
  A. The gravedigger has the only key.
  B. But he is nowhere to be seen.
  C. You can’t unlock the gate. To
     leave, go out the open gate at
     Creepy Corner.
A.1.2 Autosave
TextQuests environment automatically saves the full game state after each command. An agent can then restore to any of these previous states by issuing the special command restore {step_id}, where restore {step_id} corresponds to the step it wishes to return to. This functionality is crucial as it allows an agent to explore different strategies or recover from non-optimal paths, such as getting stuck or entering an unwinnable state, without requiring a full restart of the game. This emulates the save/restore patterns common among human players to facilitate exploration and experimentation. As shown in Figure 6, providing agents with this capability resulted in a notable improvement in overall gameplay performance.
Figure 6: Adding an AutoSave mechanism to the game environment improves the agent’s exploration efficiency.
Left: An example of evaluated LLMs makes use of the autosave and restore features to experiment with different approaches to solve an in-game puzzle. Right: As LLMs’ capabilities increase, the performance difference between runs with and without the Autosave feature widens, leading to a difference of more than 10% after 500 steps on Gemini 2.5 Pro and Claude Sonnet 4.0 and 6% on Grok 3 Mini.
Appendix B Beyond 500 Run Steps
We extended our evaluation to $800$ steps for several models and observed that the increase in game progress was minimal after the 500-step mark. This trend is illustrated in Figure˜7, with detailed metrics provided in Table˜3.
Figure 7: Game progress trajectories on TextQuests for selected frontier LLMs evaluated up to 800800 steps. The visualization shows game progress saturating after approximately 500500 steps. For detailed metrics, see Table˜3.
No Clues With Clues
500 steps 800 steps 500 steps 800 steps
Progress (%) $\uparrow$ Harm $\downarrow$ Progress (%) $\uparrow$ Harm $\downarrow$ Progress (%) $\uparrow$ Harm $\downarrow$ Progress (%) $\uparrow$ Harm $\downarrow$
[Uncaptioned image] GPT-5 37.8 17.2 41.2 26.2 71.2 21.8 75.1 34.7
[Uncaptioned image] o3 30.9 18.7 35.7 25.4 60.4 17.2 63.2 26.1
[Uncaptioned image] Gemini 2.5 Pro 23.2 15.9 25.4 39.8 60.6 25.6 63.8 36.4
[Uncaptioned image] Claude Sonnet 4 24.7 16.0 27.4 26.9 57.2 18.4 58.6 27.6
[Uncaptioned image] GPT-5-mini 15.9 12.0 21.1 17.6 42.1 15.7 45.1 21.3
Table 3: Game progress and harm for several LLMs at 500500 and 800800 run steps.
Appendix C Full Results
No Clues With Clues
Progress (%) $\uparrow$ # Completed (/25) $\uparrow$ Harm $\downarrow$ Progress (%) $\uparrow$ # Completed (/25) $\uparrow$ Harm $\downarrow$
[Uncaptioned image] GPT-5 37.8 0 17.2 71.2 5 21.8
[Uncaptioned image] Claude Opus 4.1 33.9 0 19.1 68.0 4 22.1
[Uncaptioned image] Grok 4 31.2 0 30.4 61.4 3 31.4
[Uncaptioned image] o3 30.9 0 18.7 60.4 3 17.2
[Uncaptioned image] Claude Opus 4 26.4 0 16.5 60.5 4 19.2
[Uncaptioned image] Gemini 2.5 Pro 23.2 0 15.9 60.6 3 25.6
[Uncaptioned image] Claude Sonnet 4 24.7 0 16.0 57.2 2 18.4
[Uncaptioned image] Grok 3 18.9 0 15.4 41.9 2 21.2
[Uncaptioned image] GPT-5 mini 15.9 0 12.0 42.1 1 15.7
[Uncaptioned image] GPT-4.1 22.8 0 11.4 37.5 0 15.3
[Uncaptioned image] Grok 3 mini 22.4 0 17.8 32.2 0 18.2
[Uncaptioned image] Qwen 3 Thinking 15.1 0 16.4 29.8 1 10.8
[Uncaptioned image] Gemini 2.5 Flash 14.4 0 11.7 31.8 0 16.8
[Uncaptioned image] DeepSeek R1 15.2 0 15.4 23.8 0 23.0
[Uncaptioned image] o4-mini 12.8 0 18.6 20.6 0 20.0
[Uncaptioned image] Kimi K2 10.5 0 8.3 19.7 0 9.0
[Uncaptioned image] GPT-OSS 120B 12.0 0 21.2 18.1 0 12.9
[Uncaptioned image] Gemini 2.5 Flash-Lite 11.7 0 22.8 16.6 0 10.6
[Uncaptioned image] GPT-4.1-mini 10.6 0 11.7 15.9 0 12.2
[Uncaptioned image] Claude Haiku 3.5 12.3 0 14.9 13.4 0 10.6
[Uncaptioned image] Llama 4 Maverick 9.2 0 13.1 16.1 0 12.1
[Uncaptioned image] Llama 4 Scout 4.8 0 7.4 7.7 0 7.0
Table 4: LLMs performance on TextQuests. Reasoning models are evaluated with high reasoning setting.
Appendix D System Prompt and Environment Interaction
System Prompt
We use the following system prompt with an instruction to output a structured format:
System Prompt You are a text-based game player. Your goal is to finish the game. GAME: game_name MAX SCORE: max_score During each game turn, you will be given an observation of the current game state. To take an action, generate a single parser command action. {guidelines_and_clues} [OUTPUT FORMAT DURING GAME TURNS] Return a final json block with the following format:
    "reasoning": "<your step-by-step reasoning for action>",
    "action": "<action to take>"
SPECIAL ACTION: You can restore the game to previous steps by using ‘restore {step_id}’ for the step that you want to backtrack.
Environment Inputs/Outputs
As shown in Figure˜8, the evaluation proceeds in discrete history turns. For each turn, we parse the brief reasoning and corresponding action from the LLM’s output, adding them to the context history for the subsequent step. To ensure the agent’s decisions are based on its explicit plan, we discard any other ’thinking’ output and do not include it in the context history.
Figure 8: An illustration of an agent’s turn. From the model’s output, the brief reasoning and action are extracted and added to the context history, while any intermediate ‘thinking’ is discarded
Appendix E Token Analysis
We report the total input/output tokens cost to evaluate TextQuests in Table˜5.
No Clues
o3 Gemini 2.5 Pro Claude Opus 4.0 Claude Sonnet 4.0 GPT-4.1 GPT-4.1-mini Max Input Tokens 82K 128K 140K 132K 97K 78K Max Output Tokens 6.2K 700 1.4K 1.6K 239 172 Total Input Tokens 471M 562M 524M 569M 460M 428M    Cache Tokens 450M 530M 522M 567M 456M 420M Total Output Tokens 10M 2.7M 3.1M 3.3M 0.7M 0.7M
With Clues
o3 Gemini 2.5 Pro Claude Opus 4.0 Claude Sonnet 4.0 GPT-4.1 GPT-4.1-mini Max Input Tokens 90K 132K 140K 132K 88K 97K Max Output Tokens 6.8K 1.4K 1.7K 1.9K 217 199 Total Input Tokens 531M 675M 585M 569M 509M 539M    Cache Tokens 514M 635M 583M 567M 503M 530M Total Output Tokens 9.6M 2.2M 2.8M 3.3M 0.7M 0.7M
Table 5: Input and output token costs for evaluating TextQuests. All models were configured for high reasoning effort (and a 20k token thinking budget for Claude 4 models), though this maximum budget was not always fully utilized. While the majority of the cost is from input tokens, a high cache hit rate (exceeding 95-99%) makes the evaluations significantly cost efficient.
Appendix F Comparing Game Progress and Game Score
As discussed in Section˜3.2, the built-in scoring systems of the Infocom games are often a weak proxy for an agent’s actual advancement toward completing a game. They were designed to reward human players for exploration and cleverness, not to serve as a direct measure of progress along the critical path.
To visually illustrate this discrepancy, Figure˜9 presents a direct comparison between the traditional Game Score and our checkpoint-based Game Progress metric. The figure highlights how our metric provides a more consistent signal of an agent’s approach to completion and shows clear cases where the game’s score is decoupled from this primary objective.
Figure 9: A comparison of our Game Progress metric against the in-game Game Score. Left: The trajectory for an optimal walkthrough of a sample game shows that our Game Progress provides a more representative signal of advancement than the built-in score. Right: The final scores for games like Moonmist and Witness demonstrate that game completion (100% progress) is often independent of achieving the maximum possible game score.